# Cross-Sectional Dynamics Under Network Structure: Theory and Macroeconomic Applications

### v6

@@ Abstract
Many environments in economics involve units linked by bilateral ties. I develop an econometric framework that rationalizes the dynamics of cross-sectional variables as the innovation transmission along fixed bilateral links and that can accommodate rich patterns of how network effects of higher order accumulate over time. The proposed Network-VAR (NVAR) can be used to estimate dynamic network effects, with the network given or inferred from dynamic cross-correlations in the data. In the latter case, it also offers a dimensionality-reduction technique for modeling high-dimensional (cross-sectional) processes, owing to networks’ ability to summarize complex relations among variables (units) by relatively few bilateral links. In a first application, I show that sectoral output growth in an RBC economy with lagged input-output conversion follows an NVAR. I characterize impulse-responses to TFP shocks, and I estimate that the lagged transmission of productivity shocks along supply chains can account for 61% of aggregate output growth persistence. The remainder is due to persistence in the aggregate TFP process, leaving a negligible role for persistence in sectoral TFP. In a second application, I forecast macroeconomic aggregates across OECD countries by assuming and estimating a network that underlies the dynamics. In line with an equivalence result I provide, this reduces out-of-sample mean squared errors relative to a dynamic factor model. The reductions range from -12% for quarterly real GDP growth to -68% for monthly CPI inflation.
@@ Introduction
Numerous economic environments feature a cross-section of units linked by a network of bilateral ties, such as sectors connected through supply chains or individuals acquainted to each other. A large theoretical and empirical literature [^1] See references in the following paragraph and subsequent literature review. documents that networks amplify idiosyncratic shocks and generate comovement in cross-sectional variables at a given point in time. For example, the correlation between two sectors’ production is higher the closer the two sectors are in the supply chain network, and sectoral disturbances have a stronger effect on aggregates the better connected the sector is.
What is less well understood, however, is how network-induced comovements play out over time. With regard to the timing of network effects, the literature mostly considers two special cases. The first assumes that innovations transmit via bilateral links contemporaneously (e.g. Acemoglu et al. (2012); Elliott et al. (2014)), which leads to a static framework and implies that connections of all order play out simultaneously. The second case posits that network effects materialize exactly one link per period (e.g. Long and Plosser (1983); Golub and Jackson (2010)). These assumptions are useful for theoretical insights or for empirical work involving steady state comparisons – i.e. long-term effects of permanent innovations [^2] Under one period-lagged network interactions, these are equal to the effects in the static framework of contemporaneous linkages. See Section 2.3. –, but they are of limited use for empirical studies concerned with transition dynamics.
I build an econometric framework which relates the dynamics of cross-sectional variables to the innovation transmission along bilateral links between cross-sectional units. Assuming uni-directional transmission333i.e. either downstream or upstream, the distinction being relevant only for directed networks. and time-invariant links, the model can accommodate rich patterns on how innovations travel along bilateral links over time, and, consequently, how network connections of higher order accumulate as time progresses. The interdependence of observations $y_{it}$ and $y_{j,t-h}$ arises as the interplay of the temporal distance between periods $t$ and tht-h and the cross-sectional distance between units $i$ and $j$ encoded by the network. Correspondingly, stationarity can be characterized in terms of eigenvalues of the network adjacency matrix and roots of an AR process defined by the timing of innovation transmission along a single link. This timing is fundamentally related to the frequency of network interactions relative to the frequency of observation.
@@
The Network-VAR (NVAR) is useful in two rather distinct lines of empirical work with cross-sectional time series. By conducting inference on the timing of innovation transmission along bilateral links, it can be used to estimate dynamic network effects. Thereby, the network can be taken as given or inferred from dynamic cross-correlations in the data, possibly aided by shrinking towards observed links. With both the network and effect-timing estimated, the NVAR is also applicable as a dimensionality-reduction technique for modeling high-dimensional processes. It assumes that dynamics are generated by innovation transmission along a (small) set of bilateral links among variables. Under a given network, inference on the timing of network effects boils down to a linear regression with covariates that summarize lagged observations using bilateral links. Joint inference is implemented by iterating on analytically available conditional estimators, with Bayesian as well as frequentist interpretations. I illustrate each of these two model uses with a respective application.
@@
In the first application, I show that the NVAR approximates the process of sectoral output and prices in a Real Business Cycle (RBC) input-output economy with time lags between the production of goods and their subsequent use as intermediaries in producing other goods. In turn, I evaluate the potential of such lagged input-output conversion (IOC) to generate endogenous business cycles, as argued-for theoretically by Long and Plosser (1983). For this purpose, I generalize the one period-lagged IOC in Long and Plosser (1983) by assuming that inputs from the past $p$ periods can be used in production and are imperfect substitutes – a stand-in for frictions like storage capacity constraints. Assuming a difference-stationary TFP process with persistent aggregate and sectoral shocks, I characterize impulse-responses and the differing implications of lagged IOC and persistence in exogenous shocks as sources of persistence in aggregate output growth. Using input-output tables and monthly output growth across 23 manufacturing sectors in the US economy, I estimate their respective contributions.
The results suggest that lagged IOC alone explain 61% of the autocorrelation in aggregate output growth. Together with persistence in aggregate TFP, this number grows to 93%, leaving little room for persistence in sectoral TFP processes. In addition, the analysis reveals interesting patterns on how supply chain linkages of different order map into impulse-responses. While existing literature documented that the effect of a sector’s TFP shock on other sectors’ output and – therefore – aggregate output depends on the former’s role as a direct and further upstream supplier to other sectors in the economy, my analysis under lagged IOC allows me to investigate how such supply chain linkages of different order translate into different timings of IRFs. I find no clear relationship between the strength and timing of a sector’s effect on aggregates.
With the second application I show that the NVAR – with the network itself estimated along with the timing of network effects – offers a parsimonious, yet flexible way of approximating high-dimensional (cross-sectional) processes. Parsimony is obtained because dynamics are driven by bilateral links and because variables (units) can be connected even in absence of a direct link between them. As a result, the dynamic comovement of all series can be modeled with relatively few non-zero bilateral links. This is reminiscent of the assumption that longer-term dynamics are driven by a set of shorter-term dynamics, which is upheld by the general class of VARMA($p,q$) models. Flexibility is owed to the fact that the network is estimated and that the model can accommodate fairly general patterns of how network effects of higher order accrue over time.
@@
I derive an equivalence result, which suggests that the NVAR is preferred to a factor model if cross-sectional dynamics are composed of many micro links rather than driven by a few influential units. This corresponds to the case of numerous sparse factors with differing sets of non-zero loadings across units, or, equivalently, a sparse, yet high-rank network adjacency matrix. This is corroborated with applications forecasting monthly industrial production growth, monthly CPI inflation and quarterly GDP growth across OECD countries. Estimating an underlying network as relevant for dynamics by shrinking links to zero, I obtain reductions in out-of-sample mean squared errors of 13%-68% relative to the ex-post best-performing dynamic factor model.
@@
This paper adds to the growing literature on networks in econometrics. [^4] See Bramoullé et al. (2016) and Graham (2020) for general references on networks in economics and econometrics. In particular, there is a large literature on spatial autoregressive (SAR) models. It is mostly concerned with identifying network effects (and effects of other covariates) in a static framework of contemporaneous dependencies (e.g. Manski (1993); Lee (2007); Bramoullé et al. (2009)). In contrast, I analyze lagged network effects. Among other studies that do so (e.g. Knight et al. (2016); Yang and Lee (2019)), Zhu et al. (2017) cast them in an explicit time series model, generalizing the lag length and discussing stationarity and large $T$-inference. I characterize dynamics in their model as the cross-sectional innovation transmission along network connections of different order. In turn, I further generalize this mapping between network connections and dynamics, which allows me to accommodate richer patterns of how transmission materializes over time. Moreover, I discuss how to conduct inference on the network itself – not only the effect-timing –, which allows me to illustrate the model’s merits as a dimensionality-reduction technique – not only as a tool for estimating dynamic network effects.
This study bridges the gap between static SAR models and dynamic VAR models. The NVAR maps networks into VARs by restricting VAR-innovations to transmit via bilateral links among variables (cross-sectional units). This leads to rich patterns of multi-step causality, connecting my work to Dufour and Renault (1998). It contrasts with Diebold and Yilmaz (2009, 2014), who map variance decompositions of VARs into networks with the goal of understanding dynamic connectedness. [^5] Another way to represent dynamics by graphs is offered in Barigozzi and Brownlees (2018). Relative to other studies restricting time series models using networks (e.g. Pesaran et al. (2004); Barigozzi et al. (2022); Mehl et al. (2023); Caporin et al. (2023)), [^6] Barigozzi et al. (2022) extend the model of Zhu et al. (2017) along a different dimension than I do, allowing for several connection-types to influence dynamics. I focus on a single variable per cross-sectional unit and a single type of connection among units, and I entertain the assumption of innovation transmission along bilateral links. This leads to a clear relation between the model’s time series properties on the one hand and the network and timing of network effects on the other. It also generates intuitive, analytical expressions for the estimators and allows me to examine the relation to factor models. [^7] Approaches for achieving shock identification using networks are discussed in Hipp (2020) and Dahlhaus et al. (2021). Bykhovskaya (2021) builds a time series model for the evolution of the network itself.
With the first application, I address the macroeconomic literature on production networks. It mostly assumes that firms convert inputs into outputs contemporaneously. In this environment, firm- or sector-level shocks transmit contemporaneously along supply chains, generating cross-sectional comovement and amplifying an individual sector’s effect on aggregates at a given point in time (Horvath, 2000; Carvalho, 2010; Foerster et al., 2011; Acemoglu et al., 2012; Bouakez et al., 2014; Giovanni et al., 2018; Giroud and Mueller, 2019). A notable exception is Long and Plosser (1983), who study an RBC economy with a one period-lag in IOC, which renders the transmission dynamic and generates endogenous business cycles.888 Under contemporaneous interactions, network effects are static because network effects of all order play out simultaneously. Networks can only amplify existing dynamics – obtained thanks to agents’ intertemporal optimization problems in a structural model or due to persistence in shocks – but not drive dynamics themselves. See discussion in Section 2.2. Carvalho and Reischer (2021) characterize persistence as a function of network properties in this environment.999 Based on the relationship they derive, the evolution of the supply chain network of the US economy can account for changes in persistence of aggregate industrial production. They also show that this calibrated model improves upon ARMA models in forecasting aggregate industrial production growth. Building on their work, I estimate to what extent business cycless in this model can be rationalized by lagged IOC as opposed to persistence in exogenous productivity shocks. This is analogous to the decomposition of aggregate variance under contemporaneous IOC in Carvalho (2010) and Foerster et al. (2011).
With the second application of the NVAR, I address the literature on dimensionality-reduction techniques for time series modeling. By reducing the number of parameters and using shrinkage priors (regularization), the NVAR leverages both ways of addressing the large parameter problem in the Wold representation (see Geweke (1984)). Compared to shrinkage methods like Lasso (Tibshirani, 1996) or Ridge regression,101010 See Hsu et al. (2008) and Camehl (2022) for applications of Lasso in the context of VARs. it applies shrinkage to links – which in turn summarize the information in predictors at all lags – rather than to predictors themselves, leading to additional parsimony.111111 Other approaches bridging sparse and factor models are usually interested in capturing the cross-sectional correlation in the errors left after factor extraction. See e.g. Fan et al. (2021). Compared to reduced rank regression and factor models (Velu et al., 1986; Stock and Watson, 2002), it finds the linear combination that effectively summarizes the information in lagged values by relying on bilateral links among variables (cross-sectional units). I show analytically that, with the cost of restricting factor dynamics, the NVAR naturally incorporates sparse factors as locally important nodes in the network. This suggests that it can improve upon the poor forecasting performance of factor models in settings with a high dispersion in factor loadings across series (see e.g. Boivin and Ng (2006)),121212 This notably includes the case of sparse factors and units differing in the set of factors they load on. See Onatski (2012) and Freyaldenhoven (2022) for analyses of sparse factors. as confirmed by the applications I consider.
@@
The remainder of this paper is structured as follows. The model and its properties are discussed in Section 2. Section 3 treats inference. In Section 4, I study how input-output connections shape sectoral output dynamics in the US economy, taking the network as given. In Section 5, I illustrate the merits of the NVAR as a dimensionality-reduction technique for modeling high-dimensional processes. Section 6 concludes.

### v7

@@ Abstract
Many economic environments involve units linked by a network. I develop an econometric framework that derives the dynamics of cross-sectional variables from the lagged innovation transmission along bilateral links and that can accommodate general patterns of how higher-order network effects accumulate over time. The proposed NVAR rationalizes the Spatial Autoregression as the limit under an infinitely high frequency of lagged network interactions. The factor-representation of the NVAR suggests that at the cost of restricting factor dynamics, it naturally incorporates sparse factors as locally important nodes in the network. The NVAR can be used to estimate dynamic network effects. When the network is estimated as well, it also offers a dimensionality-reduction technique for modeling high-dimensional processes. In a first application, I show that sectoral output in a Real Business Cycle-economy with lagged input-output conversion follows an NVAR. In turn, I estimate that the dynamic transmission of productivity shocks along supply chains accounts for 61% of persistence in aggregate output growth, leaving minor roles for autocorrelation in exogenous productivity processes. In a second application, I forecast macroeconomic aggregates across OECD countries by estimating a network behind global business cycle dynamics. This reduces out-of-sample mean squared errors for one-step ahead forecasts relative to a dynamic factor model by -12% (quarterly real GDP growth) to -68% (monthly CPI inflation). 11todo: 1go over all eqs; numbering required?22todo: 2REMEMBER: DISABLE TODONOTES & DELETE ToC BEFORE SHARING
@@ Introduction
In economics, we often encounter a cross-section of units linked by a network of bilateral ties, such as sectors connected through supply chains or individuals acquainted to each other. A large theoretical and empirical literature documents that networks amplify idiosyncratic shocks and generate comovement in cross-sectional variables at a given point in time. [^1] See work surveyed in Carvalho and Tahbaz-Salehi (2019); Bramoullé et al. (2016) and the following literature review.
How network-induced comovements unfold over time, however, is less well understood. The literature predominantly assumes contemporaneous network interactions (e.g. Bramoullé et al. (2009); Acemoglu et al. (2012); Elliott et al. (2014)), embodied by the Spatial Autoregressive (SAR) model. This implies that an innovation contemporaneously transmits along network connections of all orders. The resulting static framework is useful for “networked” steady state comparisons. Among the few studies that consider lagged network interactions, most posit that network effects materialize exactly one link per period (e.g. Long and Plosser (1983); Golub and Jackson (2010); Elhorst (2012)). While useful for analyzing the qualitative implications of the laggedness of network interactions, this assumption is of limited use for empirical work concerned with “networked” transition dynamics. Zhu et al. (2017) generalize the lag length, but the generality of their empirical framework and its relation to structurally motivated models of contemporaneous or single-lagged network interactions remain unclear.
I construct an econometric framework that derives the dynamics of cross-sectional variables from the lagged innovation transmission along bilateral links between cross-sectional units: the Network-VAR (NVAR). Like existing studies, I assume uni-directional transmission222i.e. either downstream or upstream, the distinction being relevant only for directed networks. and time-invariant links. Unlike existing studies, the model can accommodate general patterns on how innovations travel along links over time and, consequently, how network connections of higher order accumulate as time progresses. To obtain this generality, I distinguish between the frequency of network interactions and the frequency of observation. As the former diverges to infinity, the impulse-response to contemporaneous high-frequency disturbances and the networked covariance among observations approach their counterparts from the SAR model. The same impulse-response is also obtained when considering long-term effects of permanent innovations. By taking a stance on the timing of network effects, the NVAR goes beyond such steady state comparisons and characterizes transition dynamics.
In the NVAR, the interdependence of observations $y_{it}$ and $y_{j,t-h}$ arises as the interplay of the temporal distance between periods $t$ and tht-h and the cross-sectional distance between units $i$ and $j$ encoded by the network. Under a sparse network, this dependence is modeled parsimoniously; the dynamic comovement among all series is rationalized by the dynamic innovation transmission along a few bilateral links among units. This is akin to the assumption that longer-term dynamics are driven by a set of shorter-term dynamics, upheld by the general class of VARMA($p,q$) models. The dynamics under the NVAR can be represented by a Dynamic Factor Model (DFM) whereby the number of factors equals the rank of the network adjacency matrix. At the cost of restricting their dynamics, the NVAR naturally incorporates sparse factors as locally important nodes in the network.
@@
The NVAR is useful in two distinct kinds of work with cross-sectional time series. By conducting inference on the timing of innovation transmission along bilateral links, it can be used to estimate dynamic network effects. Thereby, the network can be taken as given or it can be inferred from dynamic cross-correlations in the data, possibly aided by shrinking towards observed links. When both the network and effect-timing are estimated, the NVAR is also applicable as a dimensionality-reduction technique for modeling high-dimensional processes. Conditional on a network, inference on the timing of network effects boils down to a linear regression with covariates that summarize lagged observations using bilateral links. Joint inference is implemented by iterating on analytically available conditional estimators, with Bayesian as well as frequentist interpretations. I illustrate each of these two model uses with a respective application.
@@
In the first application, I show that the NVAR approximates the process of sectoral output in a Real Business Cycle (RBC) input-output economy with lagged input-output conversion (IOC). In turn, I evaluate the potential of lagged IOC to generate endogenous business cycles, as suggested theoretically by Long and Plosser (1983). I generalize their one period-lagged IOC by assuming that inputs from the past $p$ periods are imperfectly substitutable in the production at any time $t$ – a stand-in for a range of frictions that prevent just-in-time input-sourcing – and by allowing the model frequency to differ from the observational frequency. I characterize impulse-responses under a difference-stationary TFP process with persistent aggregate and sectoral shocks. The analysis illustrates the differing implications of lagged IOC and persistence in exogenous shocks for the dynamics of sectoral and aggregate output; the former formalizes the idea that idiosyncratic changes in, say, energy production take time to feed through the networked economy, while the latter merely prolongs the effects of each round of transmission. Using input-output tables and monthly output growth across 23 manufacturing sectors in the US economy, I estimate their respective contributions to persistence in aggregate output growth.33todo: 3to the autocorrelation function of …?
The results suggest that lagged IOC explain 61% of the autocorrelation in aggregate output growth. With persistence in aggregate TFP, this number grows to 93%, leaving little room for persistence in idiosyncratic sectoral TFP. My analysis reveals that shocks to more central sectors not only affect aggregate output more strongly in the long run – as established by Acemoglu et al. (2012)44todo: 4CITE MORE? –, but they also tend to materialize more sluggishly. This relationship, however, is far from clear-cut; for example, owing to its position at the top of supply chains, a TFP shock in the primary metals sector materializes more slowly than a shock to the food and beveredge sector, despite similar long-term effects.55todo: 5relate results to Foerster et al. and Krishna Kamepalli et al.?
In the second application, I investigate the merits of the NVAR as a dimensionality-reduction technique for forecasting high-dimensional (cross-sectional) processes. I consider monthly industrial production growth, monthly CPI inflation and quarterly GDP growth across OECD countries. Under a sparse network, the NVAR models the dynamics parsimoniously; the dynamic comovement among all series is rationalized by the dynamic innovation transmission along a small set of bilateral links among countries. In line with that, I estimate the network by shrinking links to zero using a Lasso-penalty. I also consider alternative specifications that apply a Ridge-penalty and shrink to observable links.66todo: 6LEFT TO DO
@@
In my setting, the NVAR reduces out-of-sample mean squared errors by 13%-68% relative to the ex-post best-performing DFM. This corroborates my equivalence result: the NVAR better forecasts these series, whose cross-country dynamics seem to be driven by many micro links rather than a few influential countries. This corresponds to a setting with numerous sparse factors and differing sets of non-zero loadings across units or, equivalently, a sparse, yet high-rank network adjacency matrix. 77todo: 7add sentence on sparseness and interpretability of network I estimate
@@
There is a large econometric literature on spatial and network-interactions (Manski, 1993; Kelejian and Prucha, 2010; Anselin, 2003; Lee, 2007; Bramoullé et al., 2009; Shalizi and Thomas, 2011; Kuersteiner and Prucha, 2020; de Paula et al., 2024). It is concerned with identifying spatial/peer-effects and accounting for heterogeneity in cross-sectional and panel data-regressions. [^3] A separate strand devises methods for dyadic data, i.e. the case when edges rather than nodes in the network are observed (Graham, 2020). It predominantly assumes contemporaneous network interactions, as exemplified by the canonical SAR model. Among the few studies that consider lagged interactions (Elhorst, 2012; Knight et al., 2016), Zhu et al. (2017) cast them in an an explicit time series model, generalizing the lag length and discussing stationarity and large $T$-inference. I characterize Granger causality in their environment as the cross-sectional innovation transmission along network connections of different order, formalizing the idea that network effects materialize dynamically over time. In turn, I generalize this mapping between network connections and dynamics, which allows me to derive the SAR model as the limit of an underlying process driven by high-frequency lagged network interactions. [^4] Barigozzi et al. (2025); Maung (2022); Armillotta and Fokianos (2023) extend Zhu et al. (2017) along different dimensions than I do. Like Pinkse et al. (2002); Lam and Souza (2020); Qu et al. (2021); de Paula et al. (2024) do in the static environment, I discuss inference on the network itself – not only the effect-timing –, which institutes the NVAR as a dimensionality-reduction technique – not only a tool for estimating dynamic network effects – and motivates my equivalence result to factor models.
Many studies incorporate networks in time series analyses. Following Diebold and Yılmaz (2009, 2014), some researchers interpret quantities of a given time series model – typically a VAR – using network analysis (Billio et al., 2012; Anufriev and Panchenko, 2015; Chen et al., 2025). Other studies restrict time series models using networks (Carvalho and West, 2007; Chudik and Pesaran, 2011; Davis et al., 2016; Ahelegbey et al., 2016; Zhu et al., 2017; Barigozzi and Brownlees, 2018; Billio et al., 2019; Caporin et al., 2021; Martin et al., 2025). [^5] Holly and Petrella (2012); Dahlhaus et al. (2021); De Graeve and Schneider (2023); Chodorow-Reich et al. (2025) exploit networks for shock-identification in VARs. Bykhovskaya (2023) models the temporal evolution of a weighted network, which belongs to the realm of dyadic regressions (Graham, 2020) and (dynamic) network-formation rather than -effects. Among them, the Global VAR (GVAR) (Pesaran et al., 2004; Chudik and Pesaran, 2014) restricts static and dynamic interdependencies across units in a panel-VAR using exogenous network-weights – e.g. bilateral trade –, enabling the estimation of many interconnected VAR models under a weak exogeneity assumption. [^6] See Canova and Ciccarelli (2013) for an overview of panel-VAR approaches. The NVAR also belongs to this second category; it restricts innovations in a (cross-sectional) VAR to transmit along bilateral links among variables (units), which leads to rich patterns of multi-step causality as in Dufour and Renault (1998). In contrast to the GVAR and most “networked” time series approaches above, I focus on a single variable per cross-sectional unit and a single type of connections among them, and I entertain the assumption that innovations transmit along bilateral links. This allows me to characterize a range of theoretical properties – including the relation to static SAR models and DFMs –, to structurally motivate the NVAR with an RBC production economy, and to efficiently conduct inference by relying on analytical conditional estimators.
When the network is estimated parsimoniously, the NVAR also addresses the literature on dimensionality-reduction methods for time-series modeling. By reducing the number of parameters and applying shrinkage priors, it leverages both approaches available for addressing the large parameter problem in the Wold representation (Geweke, 1984). Compared to standard shrinkage methods (Litterman, 1986; Tibshirani, 1996), it applies shrinkage to links – which in turn summarize the information in predictors at all lags – rather than to predictors themselves. Compared to reduced rank regression (Velu et al., 1986) or factor models (Geweke, 1977; Stock and Watson, 2002), it finds the linear combination to summarize the information in predictors by relying on network connections of different order and, ultimately, on bilateral links among units. While many of the network-restricted time series approaches mentioned above appear to have a reduced-rank representation – consider the GVAR’s construction of weighted averages of other countries’ variables –, the simple setup of the NVAR allows me to characterize this representation explicitly. Owing to the equivalence result that compares the NVAR to the DFM – corroborated by their relative performance forecasting the three processes I consider –, the NVAR contributes to the long-standing literature on sparse factors (Boivin and Ng, 2006; Onatski, 2012; Bai and Ng, 2019; Fan et al., 2022; Anatolyev and Mikusheva, 2022; Freyaldenhoven, 2025). Interestingly, in my application, a sparse network is preferred to a dense one, suggesting the presence of sparse rather than weak factors and running counter to the results of Giannone et al. (2021).
With the other application of the NVAR, I contribute to the series of efforts to microfound aggregate dynamics using production networks (see Horvath (1998, 2000); Dupor (1999); Shea (2002); Carvalho (2010); Foerster et al. (2011); Acemoglu et al. (2012); di Giovanni et al. (2014) and survey of Carvalho and Tahbaz-Salehi (2019)). These studies rely on contemporaneous IOC – mirroring the prevalence of contemporaneous interactions in econometrics and in the related study of spatial economies (Desmet and Parro, 2025) – and show that the production network greatly contributes to aggregate volatility. A notable exception is Long and Plosser (1983), who show that lagged IOC endogenizes business cycles – i.e. persistence in aggregates –, without relying on autocorrelation in exogenous shocks, non-rational expectations, or other frictions. Their analysis points to a more prominent role for production networks than as mere amplification devices of existing dynamics; it formalizes (qualitatively) the notion that sectoral shocks take time to feed through the economy, an idea well-accepted in the trade literature (Alessandria et al., 2010; Liu and Tsyvinski, 2024; Antràs and Tubdenov, 2025). Carvalho and Reischer (2021) characterize persistence in the Long and Plosser (1983)-economy and show that its implied evolution due to observed changes in the US input-output network accounts well for empirical measures of the changing persistence of aggregate output growth. Their analysis suggests that the endogenization of business cycles through lagged IOC is not only theoretically attractive, but has empirical merit as well. I quantify the empirical salience of Long and Plosser (1983)’s hypothesis by taking an RBC economy with general lags in IOC to the data. While lagged IOC leads to the same steady state as under contemporaneous IOC, it decomposes the long-term responses to granular TFP shocks, with distinct transition dynamics from those generated by autocorrelated TFP processes. By conducting inference on the timing of IOC along with the autocorrelations and variance of exogenous TFP processes, I quantify the contribution of lagged IOC to the entire autocorrelation function of aggregate output growth, supplementing and extending earlier decompositions of aggregate volatility (Foerster et al., 2011; di Giovanni et al., 2014; De Graeve and Schneider, 2023).
Outline
@@
The rest of this paper is structured as follows. The model and its properties are discussed in Section 2. Section 3 treats inference. In Section 4, I study how input-output connections affect output dynamics in the US economy. In Section 5, I illustrate the merits of the NVAR as a dimensionality-reduction technique. Section 6 concludes.